2. The 2nd plot is the results in runs between 11/30 and 12/04 on which we change left arm to be 20 degree. The holes grow bigger.
3. The last plot is the current one after changing to 20 degree with the new discriminator threshold adjust.
The possible reason might be from two sources: 1. threshold too high 2. loss of one of two blocks of shower signals 3. timing match between some preshower blocks and the corresponding shower blocks.
We are going to ask for an access to fix the current problem.
Figure 1
Fig_Run25769_E_Na_Eff_Left_HardFigure 2
Figure 3